Close menu

Logbook

This section contain instruments logs for the duriation of the campaign. You can group the logs either by instrument or by date.

March 18, 2017

PollyXT lidar – PollyXT arrival and installation

1200: RE, MK. PollyXT container is now installed in Finokalia station. Notes:
i) One A/C in operation due to malfunction of the vacuum pump.
ii) Water cooler is placed inside the container.
iii) Flashlamp counts set to zero.
iv) Air bubbles in the system cleared.
v) Pre-lasing test performed. Infrared light visible on the special paper at 1700 Volts.
vi) Laser beam alignment.
vii) Internet connection with the existing cellular network of Finokalia.
viii) 1064 off until it reaches the correct temperature.

March 19, 2017

PollyXT lidar – 1064 channel activated

0845: mk. Restart polly.exe.
1232: mk. Plugged the 1064 cable. Restart polly.exe.
1242: mk. Beam down.
1318: mk. Internet is now provided by prepay vodafone (+306951124583). Compatibility with router and switch operation ould not be achieved. LTE modem directly plugged into PC.

April 1, 2017

RPG HATPRO Microwave Radiometer – Radiometer installation

radiometer installation finalized.

April 2, 2017

Scanning polarization lidar – Vertical measurements started

First vertical measurements were started during the evening. Measurements stopped due to a power failure.

April 3, 2017

Scanning polarization lidar – Vertical measurements continued

First set of vertical measurements collected. Comparissons with the Polly XT shows good agreement.

CIMEL sunphotometer – CIMEL sunphotometer installation

CIMEL #181 sunphotometer is installed and registered as "FINO" AERONET station

April 4, 2017

RPG HATPRO Microwave Radiometer – Instrument issues

The instrument stopped working during the night. Data aquisition restarted at noon

PSR sunphotometer – installation

installed. it is cloudy so no aligment could be checked so it wasn't turned on. 

PR

PREDE POM-01 sun-sky radiometer – Installation

Instrument installed

Scanning polarization lidar – Vertical measurements

QA tests performed for alignment, triggerdelay and depolarization

Test measurements performed during evening. The data looks good.

 

April 5, 2017

CIMEL sunphotometer – CIMEL lunar/sun-photometer installation

CIMEL #989 (lunar/sun-photometer) is installed and registered as "FINO" AERONET station

CIMEL #181 still measures, but it is not registered in AERONET

PSR sunphotometer – alignment

8.32 UTC The instrument was aligned 4q 

the new offsets are

horizontal -0.047

vertical -0.092

 

11.00 UTC the instrument's interface was stucked for a couple of hours. it was restarted and functioned well

 

13 UTC optical parts cleaned

 

 

 

CMP - UVB - Pyranometers – installation

13 utc 

installing both pyranometers is finished. some electronic parts are missing for correct connection to the logger. so tommorow the logger will be set up

PREDE POM-01 sun-sky radiometer – Prede POM allignement

The Prede POM was alligned at about 6:50 UTC

RPG HATPRO Microwave Radiometer – Instrument issues

The instrument stopped working during the night. Data aquisition restarted at noon. Measurement definition file changed to solve the instrument freezing issue.

Scanning polarization lidar – Vertical measurements

Vertical measurements performed. Low clowds.

April 6, 2017

CMP - UVB - Pyranometers – soft installation

13 UTC the instruments are up and running

 

12 utc cleaning domes

 also today the logger software was installed

PSR sunphotometer – cleaning

12:00 utc

optical parts cleaned

 

software was stopped for an hour and a half for installing on the laptop pyranometer software

 

 

April 8, 2017

PSR sunphotometer – Optical parts cleaning

at ~12:20 UTC

CIMEL sunphotometer – Spider

A spider is removed at ~12:30 UTC

PREDE POM-01 sun-sky radiometer – Spiders inside PREDE POM

Spider was removed from the tube of the instrument, at about 2:00 UTC

CMP - UVB - Pyranometers – Optical parts cleaning

At ~12:20 UTC

April 9, 2017

PSR sunphotometer – Power supply

8:45utc

tracker loose power supply and stuck in a position towards the east.

the problem was solved and the tracker is pointing correctly again

global measurements weren't saved till today. problem solved

 

Scanning polarization lidar – Scanning

First scans performed for three angles. 5, 30 and 60 degrees.

Multiple tests from 0 to 5 degrees performed.

April 10, 2017

CMP - UVB - Pyranometers – Optical parts cleaning

7 UTC: optical parts cleaning

PREDE POM-01 sun-sky radiometer – Optical parts cleaning

7 UTC: optical parts cleaning

PSR sunphotometer – Optical parts cleaning

7 UTC: the optical parts are cleaned

CIMEL sunphotometer – Spider behind the optical head

At ~7 UTC a spider was found behind the optical head of #181 (not inside the collimator unit) and was removed.

Scanning polarization lidar – Scanning

Measurements performed. Complex scanning scheme performed at 3 (three) angles (5, 30, 60 degrees) for N, S, E, W

April 11, 2017

CIMEL sunphotometer – clouds

Almost overcast conditions

April 12, 2017

CIMEL sunphotometer – clouds

Overcast conditions

RPG HATPRO Microwave Radiometer – Power related issues

the instrument stopped recording due to a power failure at the station. Measurements restarted.

April 13, 2017

CIMEL sunphotometer – Clouds

Overcast skies today

Scanning polarization lidar – Scanning measurements - preparation

The frame and motor for the scanning were installed. After a first set of vertical measurements, the lidar was installed on the scanning module.

April 14, 2017

PSR sunphotometer – cleaning

6-7 UTC optical parts cleaning

PREDE POM-01 sun-sky radiometer – cleaning malfunction

6-7 UTC

optical parts cleaning

during the optical parts cleaning and spider inspection a plastic ring from incide the collimator tube fell from the roof. We found it and we will put it back in place during a cloudy day

April 17, 2017

PREDE POM-01 sun-sky radiometer – cleaning

9 UTC optical parts cleaning

CIMEL sunphotometer – Inspection for spiders

We found a spider web on #989, but not any spider

PSR sunphotometer – cleaning

9 UTC optical parts cleaning

also, until 9 UTC the instrument did not work, because a cable was detached

April 18, 2017

CMP - UVB - Pyranometers – Optical parts cleaning

15:00 UTC optical parts cleaning

PSR sunphotometer – Optical parts cleaning

15:00 UTC optical parts cleaning

PREDE POM-01 sun-sky radiometer – Optical parts cleaning

15:00 UTC optical parts cleaning

April 19, 2017

CMP - UVB - Pyranometers – optical parts cleaning

16:00 UTC optical parts cleaning

PSR sunphotometer – optical parts cleaning

16:00 UTC optical parts cleaning

PREDE POM-01 sun-sky radiometer – optical parts cleaning

16:00 UTC optical parts cleaning

April 20, 2017

PSR sunphotometer – Optical parts cleaning

17:00 UTC optical parts cleaning

PREDE POM-01 sun-sky radiometer – Optical parts cleaning

17:00 UTC optical parts cleaning

CMP - UVB - Pyranometers – Optical parts cleaning

17:00 UTC optical parts cleaning

CIMEL sunphotometer – Clouds

Many clouds, almost throught the whole day

April 21, 2017

CIMEL sunphotometer – Clouds

Overcast skies today

April 22, 2017

CIMEL sunphotometer – Clouds and rain

Clouds and rain today

April 24, 2017

PSR sunphotometer – Optical parts cleaning

9:40 UTC optical parts cleaning

CMP - UVB - Pyranometers – Optical parts cleaning

9:40 optical parts cleaning

PREDE POM-01 sun-sky radiometer – Optical parts cleaning

9:40 UTC optical parts cleaning

April 27, 2017

PSR sunphotometer – Optical parts cleaning

10 UTC optical parts cleaning

CMP - UVB - Pyranometers – Optical parts cleaning

10 UTC optical parts cleaning

April 29, 2017

CIMEL sunphotometer – Clouds and spider

7 UTC cirrus clouds in the sky

 

7 UTC a spider web was removed from #989 (but no spider was found)

Contact

For further information about the campaign please contact Vassilis Amiridis.

Acknowledgement

This campaign has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 654109.

Credits

Website developed by impworks.